# Early stages of collective cell invasion: Biomechanics

### v1

@@ 2.2.2. Implementation of the Migration Force
To overcome this difficulty, we propose a fractional time step method. In this method, displacements due to stiffness-induced forces $\mathbf{f}$ are computed at time $t=n$ ($n=0,1,2,\ldots$) and Eqs. (1)-(9) are solved. After this, we find a ‘migration’ displacement $\mathbf{u}_{m}$ and the corresponding ‘migration’ strain by solving the Navier equation
@@
at the intermediate time step $t=n+1/2$. The migration displacement produces a strain
@@
$\Delta H_{mig}=-g(\mathbf{x},\mathbf{x}^{\prime})\lambda_{mig}\left[h(E(\epsilon_{m1}))(\mathbf{v}_{1}\cdot\mathbf{v}_{m})^{2}+h(E(\epsilon_{m2}))(\mathbf{v}_{2}\cdot\mathbf{v}_{m})^{2}\right].$ (11c)
We give equal weight to migration and stiffness-induced variations of energy, $\lambda_{mig}=\lambda_{strain}$. In Eq. (11c), the threshold energy $E_{\theta}$ is always that of the target pixel, both in extension and in retraction. This breaks the symmetry extension-retraction existing in the energy variation of Eq. (6) and facilitates motion of the cells in a preferred direction.
@@
  • Single time step: Combines stiffness and migration forces in a single equation ($\mathbf{K}\mathbf{u}=\mathbf{f}+\mathbf{f}_{m}$) at the same time step. While simple to implement, this approach may fail to generate effective migration displacement under certain conditions.
@@
  • Fractional time step: Separates the stiffness ($\mathbf{K}\mathbf{u}=\mathbf{f}$) and migration ($\mathbf{K}\mathbf{u}_{m}=\mathbf{f}_{m}$) force computations into distinct time steps that have distinct strain tensors hence and energy changes. This method allows iterative updates to the Hamiltonian for stiffness forces at $t=n$ and migration forces at $t=n+1/2$.
The flowcharts of the two approaches are presented in Appendix A.
@@
Fig 4: Comparison of migration force implementations. A total of $N_{total}=111$ cells are randomly placed inside a circular enclosure, with $N_{m}=19$ M cells (blue) and passive E cells (red). Parameter configuration as in Fig. 3(f): $E_{\theta c}=1$ kPa, $E_{\theta m}=15$ kPa, $j_{cs}=1$, $j_{cc}=1$, $j_{cm}=4$, $j_{ms}=1/4$, and $j_{mm}=6$. (a) Initial configuration. At $t=3000$: (b) No migration force, (c) migration forces with the single time step approach, and (d) migration forces with the fractional time step approach. The attraction point is located in the bottom-left corner.
@@
As shown in Fig 4, the fractional time step method ensures that active cells reach the attraction point away from the reference point even under challenging adhesion configurations, outperforming the single time step approach. Consider the same configuration as in Fig. 3(f). In the absence of the migration force, active cells escape from the aggregate without a preferred direction; see Fig. 4(b). What happens when we add reference/attraction points and a migration force as in Figs. 4(c) and 4(d)? Cluster migration is easier to achieve as neighboring pixels with different stiffness thresholds generate a gradient directing movement toward the attraction point. However, migration sometimes fails when traction and migration forces are combined in a single time step as in Fig. 4(c) but the fractional step approach succeeds as shown by Fig.4(d).
@@
Fig 5: Forces and strains produced by the single and fractional step methods. A cluster formed by 12 passive E cells (red) with 3 M cells (blue) on its boundary has evolved 500 Monte Carlo steps either by the fractional step method, Panels (a) and (f)-(j), or by the single step method, Panels (b)-(e). For Panel (b) (single step), (c) stiffness force $f$, (d) migration force $f_{m}$, (e) strain on ECM associated to force $f+f_{m}$. For Panel (a) (fractional step), (f) stiffness force $f$, (g) migration push force $f_{m}$, (h) strain on ECM associated to stiffness force $f$, (i) same for the migration force $f_{m}$, (j) sum of strains associated to stiffness and migration forces. Parameter values are as in Fig. 3(f) and the target point (not shown) is on the upper right corner. Panels (c)-(j) are density plots of the corresponding moduli, which are larger the darker the region is. The initial configuration is similar to that of Fig. 4(a) but with $N_{total}=15$ cells.
@@
Figure 5 shows further details on forces and strains as generated by the single step and fractional step methods on a 15-cell cluster out of which the three M cells have left. Panels (a), (f)-(j) have been produced by the fractional step method and the others by the single step method. Let us compare the strains in the ECM around the rightmost M cell (further from the reference point on the lower left corner and therefore closer to the target point on the upper right corner) as calculated by the single step method in Fig. 5(e) and the same strains calculated by the fractional step method in Fig. 5(j). The strains near the target point are greater when calculated by the fractional step method. Clearly, M cell invasion is much more facilitated by the fractional step method than by combining stiffness and migration forces as in the single step method.
Figure 6 presents two simulations of migration of M cells from a cellular aggregate formed by E and M cells with other adhesion parameters that make it harder for E and M cells to separate. The single time step approach succeeds for M cells to reach the attraction point if $j_{mm}=1/2$ (low M-M cell adhesion, cf Fig. 6(c)), but fails if $j_{mm}=1$ (larger M-M cell adhesion, cf Fig. 6(f)), whereas the fractional time step approach always enables migration; see Fig. 6(b) and 6(e).
In qualitative terms, separating the steps of stiffness traction forces and migration forces is crucial. Stiffness forces as in Eq. (1) have extension-retraction symmetry by choosing the threshold energy $E_{\theta}$ corresponding to the target pixel for extension and to the source pixel for retraction. This ensures cohesive extension and retraction of the cells. Contrastingly, the migration forces of Eqs. (10) select the threshold energy of the target pixel both in extension and in retraction, which induces directional movement. Thus, the fractional time step approach provides a robust framework for modeling active migration and stiffness-driven mechanics for different adhesion values. This method with the push forces of Eq. (10b) will be adopted in subsequent simulations.
@@
Fig 6: Migration forces and aggregate dynamics. The system consists of $N_{total}=111$ cells, including $N_{m}=10$ active cells, initially forming a circular aggregate. Parameter values are $E_{\theta c}=1$ kPa, $E_{\theta m}=15$ kPa, $j_{cs}=1$, $j_{cc}=1$, $j_{cm}=10$, and $j_{ms}=1$. First row, $j_{mm}=1/2$: (a) Initial configuration at $t=15$, (b) configuration at $t=3000$ with the fractional time step approach differentiating between mechanical (stiffening) and active (migration) forces, and (c) configuration at $t=3000$ with the single time step approach. Second row, $j_{mm}=1$: (d) Initial configuration at $t=15$, (e) configuration at $t=3000$ with the fractional time step approach, and (f) configuration at $t=3000$ with the single time step approach. The point of attraction is located in the upper-right corner of the enclosure. Cells in red have passive E phenotype, while cells in blue have active M phenotype. The full dynamics corresponding to panel (b) is in the Video S3.
Computational Time Analysis.
@@
Besides the qualitative differences in outcome explained above, the computational cost of the single step and fractional step methods varies significantly. We shall compare the results of the fractional step approach in Figs. 4(d), 6(b) and 6(e) with those of the single step approach in Figs. 4(c), 6(c) and 6(f). Times were measured using a HP Laptop 15s-fq5xxx with processor 12th Gen Intel® Core™  i7-1255U (12 cores at 1.7 GHz), 16 GB RAM and 64 bit Windows 11 Home OS.
@@
  • Fig. 4(d) (fractional time step method) requires approximately 2.5 hours, compared to 45 minutes for Fig. 4(c) (single time step method).
@@
  • Fig. 6(f) (single time step) is completed in 1 hour, while Fig. 6(e) (fractional time step) takes 2 hours and 15 minutes.
@@
  • For Fig. 6(c) (single time step), the runtime is 1 hour and 45 minutes, compared to 2 hours and 15 minutes for Fig. 6(b) (fractional time step).
@@
When migration fails under the single time step approach, Fig. 6(f), the fractional time step method requires more than double the computational time (over 100% longer, Fig. 6(e)). However, in cases where migration from the aggregate succeeds under single time step, as in Fig. 6(c), the additional computational cost for the fractional time step method of Fig. 6(b) is modest, approximately 30 minutes, representing a 30% increase ($t_{f}=1.3t_{s}$, where $t_{f}$ and $t_{s}$ denote runtimes for the fractional and single time step methods, respectively). With respect to the single time step method, the computational cost of the fractional time step method does not scale linearly with its increased complexity. Precisely, the computational calculation is 2 times longer but the computational time is only 1.3 times larger. Thus, the computational cost is larger for the fractional step method and there are instances that fractional step captures the escape of individual cells while the single step method does not. In conclusion, while the fractional time step method introduces a slight additional computational cost, its robustness in enabling migration under varied adhesion parameters justifies its application in these scenarios.
3. Results
@@
Fig 7: Computing the Migration Force with Clusters. A total of $N_{total}=111$ cells are randomly placed inside a circular enclosure, with $N_{m}=10$ active cells (blue) and passive adhesive cells (red). Parameter configuration: $E_{\theta c}=1$ kPa, $E_{\theta m}=15$ kPa, $j_{cs}=1$, $j_{cc}=1$, $j_{cm}=10$, $j_{ms}=1$, and $j_{mm}=1/2$. Snapshots at: (a) $t=15$, (b) $t=300$, (c) $t=2700$, and (d) $t=3000$ with parallel migration force. The attraction point is located in the upper-right corner. See Supplementary Video S3.
@@
Fig. 7 shows a circular aggregate of 111 cells, ten of which have M phenotype and are randomly placed, and the rest are E cells. We observe intrinsic patterns in Figs. 7(a) and 7(b), such as the elongation of individually migrating cells at the aggregate outer rim, surrounded by round E cells. The M cells escape and migrate effectively toward the attracting point at the upper right corner. Figs. 7(b)-(d) show other patterns: clusters of two M cells that move forward, break bonds with E cells and deform the outer rim of the aggregate; see Supplementary Video S3. Isolated M cells surrounded by E cells are trapped and cannot escape.
@@
Hybrid E/M cells form clusters, migrate collectively. These cells are more adhesive than M cells but retain active traits, making them effective in metastasis. To model them, we assign them the same energy threshold as M cells, $E_{\theta m}=15$ kPa, whereas E cells have $E_{\theta c}=1$ kPa. In addition, we give hybrid cells specific adhesion parameters: $j_{hs}$, $j_{hc}$, $j_{hm}$, $j_{hh}$ (hybrid-substrate, hybrid-passive, hybrid-active, and hybrid-hybrid adhesion, respectively).
@@
Fig 8: Simulation of adhesive, active, and hybrid cells. A total of $N_{total}=111$ cells are randomly placed inside a circular enclosure, with $N_{m}=3$ active cells (blue), $N_{h}=4$ hybrid cells (light blue), and passive adhesive cells (red). Parameter configuration: $E_{\theta c}=1$ kPa, $E_{\theta m}=15$ kPa, $j_{cs}=1$, $j_{cc}=1$, $j_{cm}=4$, $j_{ms}=1/4$, $j_{mm}=6$, $j_{hs}=2$, $j_{hc}=10$, $j_{hm}=10$, abd $j_{hh}=1/2$. Snapshots at: (a) $t=20$. (b) $t=1500$. (c) $t=3000$. active and hybrid cells are initially positioned near the attraction point in the upper-right corner. See Supplementary Video S4.
@@
Fig 8 illustrates three cell types: E (red), M (blue), and E/M hybrid (light blue). M and hybrid cells are placed near the top corner, simulating local influences on phenotype. passive E cells have adhesion parameters that favor attachment to their own type, $j_{cs}=1$, $j_{cc}=1$, M cells have adhesion parameters that encourage them to remain isolated and interact primarily with the substrate, $j_{ms}=1/4$, $j_{cm}=4$, $j_{mm}=6$, $j_{hm}=10$, whereas hybrid cells exhibit a preference for clustering with their own kind but retain active behavior: $j_{hh}=1/2$, $j_{hs}=2$, $j_{hc}=10$. Fig. 8 shows that $N_{h}$ hybrid cells form clusters and consistently advance further than $N_{m}$ M cells within the same time frame, which is consistent with known observations fri12 ; vil21 . See Supplementary Video S4 for the complete process.
@@
Fig 9: Invasion of tissue from fingering instability of a circular aggregate of epithelial cells. Parameter configuration: $N_{total}=111$ cells, $E_{\theta c}=10$ kPa, $j_{cs}=1$, $j_{cc}=2$. Snapshots at: (a) $t=30$. (b) $t=500$. (c) $t=1000$. Note that individual cells and small cell clusters detach from the fingers. See Supplementary Video S5.
@@
A circular aggregate of epithelial cells can exhibit a fingering instability as in wound healing assays hak17 ; pet10 . This has already been captured by the CPM without migrating cells (see Figure 6 of Ref. vanOers ); see also Fig. 9. In the latter figure, the fingers of E cells resulting from the instability of the circular collectively invade the surrounding tissue. It is possible to model collective invasion of E cells led by a basal cell which is active and sense the attraction point whereas all other E cells are not active. See Fig. 10. Another possibility is to modify the adhesion parameter of a single E/M hybrid cell, which is justified as it is known that the EMT transition may exhibit different stages of completion pas18 ; bra18 ; kro19 .
@@
Fig 10: Invasion of tissue from finger formation led by a basal cell. Parameter configuration: $N_{total}=111$ cells, $E_{\theta c}=10$ kPa, $j_{cs}=1$, $j_{cc}=1/2$. Snapshots at: (a) $t=30$. (b) $t=500$. (c) $t=1000$. (d) $t=2000$. (e) $t=3000$. Note that a small cell cluster led by the green basal cell has detached from the finger. See Supplementary Video S6.
4. Discussion

### v2

@@ IV.0.2 Fractional step CPM
To enhance motility, we propose a fractional time step method: Displacements due to cell traction forces $\mathbf{f}$ are computed at time $t=n$ ($n=0,1,2,\ldots$) and Eqs. (1)-(9) solved. Then we find a ‘migration’ displacement $\mathbf{u}_{m}$ from solving
@@
at the intermediate time step $t=n+1/2$. $\mathbf{u}_{m}$ produces a strain
@@
$\Delta H_{\text{mig}}\!=-g(\mathbf{x},\mathbf{x}^{\prime})\lambda_{mig}\!\left[h(E(\epsilon_{m1}))(\mathbf{v}_{1}\cdot\mathbf{v}_{m})^{2}\!+h(E(\epsilon_{m2}))(\mathbf{v}_{2}\cdot\mathbf{v}_{m})^{2}\right]\!\!.\quad$ (11c)
We give equal weight to migration and stiffness-induced variations of energy, $\lambda_{\text{mig}}=\lambda_{\text{duro}}$. In Eq. (11c), the threshold energy $E_{\theta}$ is always that of the target pixel, both in extension and in retraction. This breaks the symmetry extension-retraction existing in the energy variation of Eq. (6) and facilitates motion of the cells in a preferred direction.
@@
  • Single time step: Combines traction and migration forces in a single equation ($\mathbf{K}\mathbf{u}=\mathbf{f}+\mathbf{f}_{m}$) at the same time step. The energy change is given by Eq. (6) with the function $h(E)$ of Eq. (9) in which the YMT may correspond to the target pixel in extension and it corresponds to the source pixel in retraction (choice A) or always to the target pixel (choice B). While simple to implement, this approach may not generate effective migration displacement as explained before.
@@
  • Fractional time step: Separates traction ($\mathbf{K}\mathbf{u}=\mathbf{f}$) and migration ($\mathbf{K}\mathbf{u}_{m}=\mathbf{f}_{m}$) force computations into distinct time steps that have distinct strain tensors and therefore different energy changes. While the durotaxis (traction force) half step keeps the extension-retraction symmetry, in the migration half step, the YMT is that of the target pixel, both in extension and in retraction. The fractional step CPM allows iterative updates to the Hamiltonian for traction forces at $t=n$ and migration forces at $t=n+1/2$.
The flowcharts of the two approaches are presented in Appendix B.
@@
Fig 4: Comparison of migration force implementations. $N_{total}=111$ cells of which $N_{m}=19$ are motile (blue) are randomly placed inside a circular enclosure. Parameter configuration as in Fig. 2(f): $E_{\theta c}=1$ kPa, $E_{\theta m}=15$ kPa, $j_{cs}=1$, $j_{cc}=1$, $j_{cm}=4$, $j_{ms}=1/4$, and $j_{mm}=6$. (a) Initial configuration. At $t=3000$ MCTS: (b) No migration force, (c) migration forces with the single time step approach, and (d) migration forces with the fractional time step approach. The attraction point is located on the top-right corner.
@@
To recapitulate, as shown in Fig 4, the fractional time step CPM ensures that motile cells reach the attraction point even under challenging adhesion configurations, outperforming the single time step approach. Consider Pattern 6 in Fig. 2(f). Without the migration force, motile cells escape from the aggregate with no preferred direction; see Fig. 4(b). What happens when we add reference/attraction points and a migration force as in Figs. 4(c) and 4(d)? Cluster migration is easier to achieve as neighboring pixels with different YMTs generate a gradient directing movement towards the attraction point. Sometimes migration does not occur with traction and migration forces combined in a single time step, cf Fig. 4(c), but the fractional step approach succeeds, cf Fig. 4(d).
@@
Fig 5: Forces and strains from the single step method. A cluster of 12 E cells (red) and 3 motile M cells (blue) on its boundary has evolved 500 MCTS by the single step method: (a) final state, and density plots of: (b) traction force $f$, (c) push migration force $f_{m}$, (d) strain on ECM associated to force $f+f_{m}$. Parameter values are as in Fig. 2(f) and the attraction point (not shown) is on the upper right corner. The initial configuration is similar to that of Fig. 4(a).
Fig 6: Forces and strains produced by the fractional step method. As in Fig. 5, (a) final state, and density plots of (b) stiffness force $f$, (c) migration push force $f_{m}$, (d) strain on ECM associated to stiffness force $f$, (e) same for the migration force $f_{m}$, (f) sum of strains associated to stiffness and migration forces. The attraction point (not shown) is on the upper right corner.
@@
Figures 5 and 6 show further details on forces and strains as generated by the single step and fractional step CPMs, respectively. Out of a 15-cell cluster, three M cells have left towards an attraction point on the upper right corner. Let us compare the strains in the ECM around the rightmost M cell (closer to the attraction point) as calculated by the single step CPM in Fig. 5(d) and the same strains calculated by the fractional step CPM in Fig. 6(f). The strains near the attraction point are greater when calculated by the fractional step CPM.
@@
Fig 7: Migration forces and invasion dynamics. Initially, $N_{m}=10$ motile blue cells ($E_{\theta m}=15$ kPa) and 101 passive red cells ($E_{\theta c}=1$ kPa) form a circular aggregate with $j_{cs}=1$, $j_{cc}=1$, $j_{cm}=10$, and $j_{ms}=1$. First row has $j_{mm}=1/2$. Configurations at MCTS: (a) $t=15$, (b) $t=3000$ (fractional step), and (c) $t=3000$ (single step). Second row has $j_{mm}=1$. Configurations at MCTS: (d) $t=15$, (e) $t=3000$ (fractional step), and (f) $t=3000$ (single step). Attraction point (not shown) located in the upper-right corner of the enclosure. The full dynamics corresponding to panel (b) is in the Video S3 suppl .
Figure 7 presents two simulations of migration of M cells from a cellular aggregate formed by E and M cells with other adhesion parameters that make it harder for E and M cells to separate. The single time step approach succeeds for M cells to reach the attraction point if $j_{mm}=1/2$ (low M-M cell adhesion, cf Fig. 7(c)), but it does not produce migration if $j_{mm}=1$ (larger M-M cell adhesion, cf Fig. 7(f)), whereas the fractional time step approach always enables migration; see Fig. 7(b) and 7(e). In qualitative terms, separating the steps of stiffness traction forces and migration forces is crucial. Stiffness forces as in Eq. (1) have extension-retraction symmetry by choosing the YMT of the target pixel for extension and that of the source pixel for retraction. This ensures cohesive extension and retraction of the cells. The selection of theYMT of the target pixel both in extension and in retraction for the migration force of Eq. (10b) enforces directional movement. Thus, the fractional CPM provides a robust framework for modeling active migration and stiffness-driven mechanics for different adhesion values and will be adopted in subsequent simulations.
IV.0.3 Computational Time Analysis.
@@
Besides the qualitative differences in outcome explained above, the computational costs of the single step and fractional step methods vary significantly. We shall compare the results of the fractional step CPM in Figs. 4(d), 7(b) and 7(e) with those of the single step CPM in Figs. 4(c), 7(c) and 7(f). Times were measured using a HP Laptop 15s-fq5xxx with processor 12th Gen Intel® Core™  i7-1255U (12 cores at 1.7 GHz), 16 GB RAM and 64 bit Windows 11 Home OS.
@@
  • Fig. 4(d) (fractional step CPM) requires approximately 2.5 hours, compared to 45 minutes for Fig. 4(c) (single step CPM).
@@
  • Fig. 7(f) (single step CPM) is completed in 1 hour, while Fig. 7(e) (fractional step CPM) takes 2 hours and 15 minutes.
@@
  • For Fig. 7(c) (single step CPM), the runtime is 1 hour and 45 minutes, compared to 2 hours and 15 minutes for Fig. 7(b) (fractional step CPM).
@@
When migration does not occur under the single step CPM, Fig. 7(f), the fractional step CPM requires more than double the computational time (over 100% longer, Fig. 7(e)). However, in cases where migration from the aggregate succeeds under single time step, as in Fig. 7(c), the additional computational cost for the fractional time step method of Fig. 7(b) is modest, approximately 30 minutes, representing a 30% increase ($t_{f}=1.3t_{s}$, where $t_{f}$ and $t_{s}$ denote runtimes for the fractional and single step CPMs, respectively). With respect to the single step CPM, the computational cost of the fractional step CPM does not scale linearly with its increased complexity. Although the operation count doubles, the runtime only increases by a factor of 1.3. In conclusion, while the fractional step CPM introduces a slight additional computational cost, its robustness in enabling migration under varied adhesion parameters justifies its application in these scenarios.
V Results
@@
Fig 8: Migration of mesenchymal cells from a cluster. Initially, $N_{m}=10$ blue M cells are randomly placed in a circular aggregate with 101 red passive cells with $E_{\theta c}=1$ kPa, $E_{\theta m}=15$ kPa, $j_{cs}=1$, $j_{cc}=1$, $j_{cm}=10$, $j_{ms}=1$, and $j_{mm}=1/2$. Snapshots at MCTS: (a) $t=15$, (b) $t=300$, (c) $t=2700$, and (d) $t=3000$. Attraction point located at the upper-right corner. See Video S3 suppl .
@@
Fig. 8 shows a circular aggregate of 111 cells, with ten randomly placed blue M cells, and 101 red E cells. We observe intrinsic patterns in Figs. 8(a) and 8(b), such as the elongation of individually migrating cells at the aggregate outer rim, surrounded by round E cells. The M cells escape and migrate effectively toward the attracting point at the upper right corner. Figs. 8(b)-(d) show other patterns: clusters of two M cells that move forward, break bonds with E cells and deform the outer rim of the aggregate; see Video S3 in suppl . Isolated M cells surrounded by E cells are trapped and cannot escape.
@@
Hybrid E/M cells form clusters, migrate collectively. These cells are more adhesive than M cells but retain motile traits, making them effective in metastasis. To model them, we assign them the same energy threshold as M cells, $E_{\theta m}=15$ kPa, whereas E cells have $E_{\theta c}=1$ kPa. In addition, we give hybrid cells specific adhesion parameters: $j_{hs}$, $j_{hc}$, $j_{hm}$, $j_{hh}$ (hybrid-substrate, hybrid-passive, hybrid-motile, and hybrid-hybrid adhesion, respectively).
@@
Fig 9: Simulation of adhesive, mesenchymal, and hybrid cells. $N_{total}=111$ cells comprising $N_{m}=3$ blue motile M cells, $N_{h}=4$ light blue hybrid E/M cells, and red passive E cells are randomly placed inside a circular enclosure, with motile and hybrid E/M cells situated closer to the attraction point at the upper-right corner. Parameter configuration: $E_{\theta c}=1$ kPa, $E_{\theta m}=15$ kPa, $j_{cs}=1$, $j_{cc}=1$, $j_{cm}=4$, $j_{ms}=1/4$, $j_{mm}=6$, $j_{hs}=2$, $j_{hc}=10$, $j_{hm}=10$, and $j_{hh}=1/2$. Snapshots at MCTS: (a) $t=20$, (b) $t=1500$, (c) $t=3000$. See Video S4 suppl .
@@
Fig 9 illustrates three cell types: E (red), M (blue), and E/M hybrid (light blue). M and hybrid E/M cells are placed near the top corner, simulating local influences on phenotype. passive E cells have adhesion parameters that favor attachment to their own type, $j_{cs}=1$, $j_{cc}=1$, M cells have adhesion parameters that encourage them to remain isolated and interact primarily with the substrate, $j_{ms}=1/4$, $j_{cm}=4$, $j_{mm}=6$, $j_{hm}=10$, whereas hybrid E/M cells exhibit a preference for clustering with their own kind but retain motile behavior: $j_{hh}=1/2$, $j_{hs}=2$, $j_{hc}=10$. Fig. 9 shows that $N_{h}$ hybrid E/M cells form clusters and consistently advance further than $N_{m}$ M cells within the same time frame, which is consistent with known observations fri12 ; vil21 . See Video S4 suppl for the complete process.
@@
Fig 10: Invasion of tissue from fingering instability of a circular aggregate of epithelial cells. Parameter configuration: $N_{total}=111$ cells, $E_{\theta c}=E_{\theta m}=10$ kPa, $j_{cs}=1$, $j_{cc}=2$. Snapshots at MCTS: (a) $t=30$. (b) $t=500$. (c) $t=1000$. Note that individual cells and small cell clusters detach from the fingers. See Video S5 in suppl .
@@
A circular aggregate of epithelial cells can exhibit a fingering instability as in wound healing assays hak17 ; pet10 . This has already been captured by the CPM without migrating cells (see Figure 6 of Ref. vanOers ); see also Fig. 10. In the latter figure, the fingers of E cells resulting from the instability of the aggregate collectively invade the surrounding tissue. It is possible to model collective invasion of E cells led by a basal cell which is motile and senses the attraction point, whereas all other E cells are not motile. See Fig. 11. Another possibility is to modify the adhesion parameter of a single E/M hybrid cell, which is justified as it is known that the EMT transition may exhibit different stages of completion pas18 ; bra18 ; kro19 .
@@
Fig 11: Invasion of tissue from finger formation led by a basal cell. Parameter configuration: $N_{total}=111$ cells, $E_{\theta c}=E_{\theta m}=10$ kPa, $j_{cs}=1$, $j_{cc}=1/2$. Snapshots at MCTS: (a) $t=30$. (b) $t=500$. (c) $t=1000$. (d) $t=2000$. (e) $t=3000$. Note that a small cell cluster led by the green basal cell has detached from the finger. See Video S6 in suppl .
VI Discussion